SB The inventory records for Radford...

[The following information applies to the questions displayed below.]

The inventory records for Radford Co. reflected the following:

Beginning inventory @ May 1 1,000 units @ $3.60
First purchase @ May 7 1,100 units @ $3.80
Second purchase @ May 17 1,300 units @ $3.90
Third purchase @ May 23 900 units @ $4.00
Sales @ May 31 3,300 units @ $5.50

TB MC Qu. 05-42 What is the amount of ending inventory assuming…

1. What is the amount of ending inventory assuming the FIFO cost flow method?

2. What is the amount of gross margin assuming the weighted-average inventory cost flow method? (Round your intermediate calculations to two decimal places.)

3. What is the amount of gross margin assuming the FIFO cost flow method?
Answer #1
FIFO
Purchase Sale Closing Inventory
Date Units Rate Total Units Rate Total Units Rate Total
01-May 1000 3.6 3600 1000 3.6 3600
07-May 1100 3.8 4180 1000 3.6 3600
1100 3.8 4180
17-May 1300 3.9 5070 1000 3.6 3600
1100 3.8 4180
1300 3.9 5070
23-May 900 4 3600 1000 3.6 3600
1100 3.8 4180
1300 3.9 5070
900 4 3600
31-May 3300 5.5 18150 100 3.9 390
900 4 3600
1. Assuming FIFO Method, Closing Inventory is (3600+390) = 3,990
3. Gross Margin using FIFO Method
Sale Value = 18,150
Cost of Sales
(1000*3.6) 3600
(1100*3.8) 4180
(1200*3.9) 4680
Purchase Cost = 12460
Gross Margin = 18,150 - 12,460 5690
2. WAM Method
Purchase Sale Closing Inventory
Date Units Rate Total Units Rate Total Units Rate Total
01-May 1000 3.6 3600 1000 3.60 3600
07-May 1100 3.8 4180 2100 3.70 7780
17-May 1300 3.9 5070 3400 3.78 12850
23-May 900 4 3600 4300 3.83 16450
31-May 3300 5.5 18150 1000 3.83 3825.6
Sales Value = 18150
Cost of Sales = (3300*3.83) 12639
Gross Margin : 18,150 - 12,639 5511
Answers
1 $ 3,990
2 $ 5,511
3 $ 5,690